# HarmonicEval: Multi-modal, Multi-task, Multi-criteria Automatic Evaluation Using a Vision Language Model

### v1

@@ Abstract
Vision-language models (VLMs) have shown impressive abilities in text and image understanding. However, existing metrics for evaluating the text generated by VLMs focus exclusively on overall quality, leading to two limitations: 1) it is challenging to identify which aspects of the text need improvement from the overall score; 2) metrics may overlook specific evaluation criteria when predicting an overall score. To address these limitations, we propose HarmonicEval, a reference-free evaluation metric that aggregates criterion-wise scores to produce the overall score in a bottom-up manner. Furthermore, we construct the Multi-task Multi-criteria Human Evaluation (MMHE) dataset, which comprises 18,000 expert human judgments across four vision-language tasks. Our experiments demonstrate that HarmonicEval achieves higher correlations with human judgments than conventional metrics while providing numerical scores for each criterion.
@@ 1 Introduction
Automatic evaluation of text generated by vision-language models (VLMs) is essential for enhancing their robustness and reliability in tasks such as image captioning and visual question answering Hessel et al. (2021); Lee et al. (2024b). As VLMs advance rapidly, the need for more informative and accurate metrics aligned with human perception has grown. However, most existing metrics focus solely on measuring the overall quality of texts generated by VLMs, lacking two critical aspects expected in automatic evaluation: 1) numerical explainability and 2) comprehensiveness.
Figure 1: Necessity of multi-criteria evaluation. (a) Example of text quality evaluation in referring expression generation. HarmonicEval provides scores for each criterion. (b) Example in visual question answering where existing interpretable approach overlooks conciseness.
First, evaluation metrics that provide only overall scores Papineni et al. (2002); Zhang et al. (2020); Hessel et al. (2021) cannot identify specific deficiencies in aspects of the text, such as clarity or fluency, as illustrated in Figure 1(a). While recent metrics for vision-language tasks Chan et al. (2023); Lee et al. (2024b) offer free-form textual explanations to complement scores, these explanations lack the granularity and consistency that come from explicit numerical breakdowns. As a result, providing a numerical explanation that links specific aspect qualities to overall scores remains a challenge in the automatic evaluation of vision-language tasks.
@@
Figure 2: HarmonicEval framework consists of three steps. (a) Criterion-wise scoring is performed by prompting a VLM to evaluate the input text based on each criterion. (b) Score smoothing improves robustness based on the first-order statistics of the output token probability distributions. (c) Score aggregation produces an overall score using harmonic weighting based on the second-order statistics.
Second, as discussed in previous studies Kasai et al. (2022); Fabbri et al. (2021), existing metrics can fail to consider specific evaluation criteria when predicting overall scores. For instance, criteria such as correctness and completeness might be prioritized while conciseness is overlooked, as illustrated in Figure 1 (b). This limitation can become more pronounced when handling a range of vision-language tasks beyond image captioning, where comprehensive evaluation is crucial to achieve better alignment with human judgments.
To address these limitations, we introduce HarmonicEval, a reference-free evaluation metric that aggregates criterion-wise scores to produce overall scores in a bottom-up manner. HarmonicEval can evaluate the quality of text across diverse vision-language tasks while providing numerical evaluation scores for each criterion, such as correctness, fluency, and conciseness. The evaluation procedure of HarmonicEval consists of three steps: 1) criterion-wise scoring, performed by prompting a VLM to evaluate text based on each specific criterion; 2) score smoothing, which enhances robustness by calculating the expected score from the output token probabilities assigned by the VLM; and 3) score aggregation, where the overall score is computed using a harmonic weighting scheme that places greater trust in criteria with lower variance, thus ensuring balanced and reliable assessments. By following these steps, HarmonicEval produces overall scores along with criterion-wise scores, offering a numerically explainable and comprehensive evaluation.
@@
Furthermore, to promote research in this field, we provide the multi-task multi-criteria human evaluation (MMHE) dataset, a new human evaluation dataset spanning four vision-language tasks and five evaluation criteria. Specifically, MMHE consists of 18,000 human expert judgments for referring expression generation (REG), visual question answering (VQA), visual document understanding (VDU), and image captioning (IC). This dataset serves as a comprehensive benchmark for evaluating the alignment between automatic evaluation metrics and human judgments, which is essential for enhancing the reliability of VLMs.
Our experiments on MMHE demonstrate that HarmonicEval exhibits higher correlations with human judgments than conventional metrics, while providing criterion-wise scores. We also show that existing metrics indeed ignore certain evaluation criteria in some cases, resulting in less comprehensive evaluations. Furthermore, we evaluate HarmonicEval on the Flickr8k-EX and Flickr8k-CF Hodosh et al. (2013) datasets and show that HarmonicEval performs comparable to or better than state-of-the-art metrics.
2 Method
This section introduces HarmonicEval, a reference-free evaluation metric that integrates criterion-wise evaluation scores through harmonic weighting. Figure 2 shows an overview of HarmonicEval. First, a VLM is prompted to generate raw evaluation scores for the input text on each criterion independently (§2.1). Second, score smoothing is applied to obtain robust scores by leveraging the output token probabilities assigned by the VLM (§2.2). Finally, the individual scores are aggregated to produce an overall score through harmonic weighting (§2.3).
2.1 Prompts for criterion-wise evaluation
@@
Table 1: Evaluation criteria.
The prompt $\bm{p}_{c}$ instructs the VLM to evaluate text with respect to the criterion $c$ on a five-point scale. Below is an example of the correctness criterion for the IC task.
@@
Here, the underlined portions indicate task-dependent variations in the content. For example, “caption” can be “answer” for the VQA task [^1] Details of evaluation and prompts are in Appendix C.1..
2.2 Score smoothing
To improve alignment with human judgment, this step applies score smoothing. Inspired by previous studies Liu et al. (2023b); Lee et al. (2024b), the first-order statistics of the output token probability distribution is utilized. Specifically, the smoothed scores $\tilde{s}_{c}$ are expected scores calculated as follows:
$\tilde{s}_{c}=\sum_{r=1}^{5}rp_{c}(r),$ (1)
where $p_{c}(r)$ is a probability of the token $r$ with respect to the criterion $c$ assigned by the VLM.
2.3 Score aggregation
Given a set of criterion-wise scores $\{\tilde{s}_{c}:c\in\mathcal{C}\}$, this step aggregates $\tilde{s}_{c}$ using harmonic weighting, which calculates the overall score $\bar{s}$ as
$\bar{s}=\sum_{c\in\mathcal{C}}w_{c}\tilde{s}_{c},$ (2)
where $0\leq w_{c}\leq 1$ are weighting coefficients. The weighting coefficients are determined based on the second-order statistics of $p_{c}$ as
$w_{c}=\frac{1}{H}\sigma_{c}^{-2(1-\gamma)/\gamma},$ (3)
where $\sigma_{c}$ is the standard deviation of the output token probabilities given by
$\sigma_{c}=\sqrt{\sum_{r=1}^{5}(r-\tilde{s}_{c})^{2}\ p_{c}(r)},$ (4)
and $H=\sum_{c}\sigma_{c}^{-2(1-\gamma)/\gamma}$ is the harmonic mean of the variances with a hyperparameter $\gamma$. Note that smaller values of $\sigma_{c}$ can be interpreted as indicating higher confidence in the evaluation of $c$.
The role of hyperparameter $\gamma$ is to bridge three weighting strategies: uniform weighting, inverse variance weighting and selective weighting. Below, we detail each strategy and discuss the hyperparameter setting.
Uniform weighting.
When $\gamma$$=$$1$, harmonic weighting reduces to uniform weighting $w_{c}$$=$$1/|\mathcal{C}|$. This is effective when all criterion-wise scores are equally reliable in determining overall scores. However, this may not be the best estimator as observed variances are ignored in the aggregation process.
Inverse variance weighting.
When $\gamma$$=$$0.5$, harmonic weighting reduces to inverse variance weighting $w_{c}\propto\sigma_{c}^{-2}$. This provides the best linear unbiased estimator under the assumption that the observed variance is due to statistical fluctuations. However, this assumption is not always reasonable, as each criterion may have its own variance.
Selective weighting.
Discussion. Considering that the evaluation criteria are carefully designed, the uniform weighting ($\gamma=1.0$) may align closely with human expert judgment, and $1.00.5\leq\gamma<$ 1.0 may further improve the alignment because it reflects the confidence of criterion-wise scores. Since the assumption underlying the inverse variance weighting ($\gamma=0.5$) may not be reasonable, we hypothesize that a value between 0.5 and 1.0 is optimal and choose $\gamma=0.75$ as the default value. The detailed mathematical background is provided in Appendix A.
3 Dataset
This section presents the MMHE dataset, which consists of 18,000 human expert judgments covering four vision-language tasks and five evaluation criteria. This dataset promotes research towards automatic evaluation that aligns with human evaluation in a fine-grained manner. Example human judgment scores are shown in Figure 3.
3.1 Vision-language tasks
To assess the multifaceted capabilities of VLMs, we selected four vision-language tasks: REG, VQA, VDU, and IC.
Referring expression generation (REG)
aims to generate textual expressions that uniquely identify a specific object marked by a bounding box. To evaluate the expression quality, the VLM is given an expression and an image with the relevant bounding box highlighted in red.
Visual question answering (VQA)
aims to answer a question posed in natural language based on the content of an image. To evaluate the answer quality, the VLM is provided with a triplet consisting of an image, a question, and an answer.
Visual document understanding (VDU)
focuses on interpreting and extracting meaningful information from visually presented documents. Quality evaluation follows a question-answering manner. Specifically, the VLM is provided with a document image, along with a question-answer pair related to the content.
Image captioning (IC)
aims to generate captions for given images. To evaluate caption quality, the VLM is provided with a text-image pair.
@@ 3.2 Dataset construction
The dataset construction process consists of three steps: 1) Source selection, selecting input texts and images for tasks from datasets; 2) Target generation, creating target texts from sources using state-of-the-art VLMs; and 3) Human expert evaluation, assessing the quality of the generated texts.
@@
We selected the following four datasets: RefCOCO Kazemzadeh et al. (2014) for REG, OK-VQA Marino et al. (2019) for VQA, VisualMRC Tanaka et al. (2021) for VDU, and MSCOCO Lin et al. (2014) for IC. We randomly sampled 100 instances from the validation or test subsets of each dataset.
@@
The target texts to be evaluated were generated using state-of-the-art VLMs. Specifically, we employed ten VLMs: LLaVA-1.5-7B/13B Liu et al. (2023a), InstructBLIP-Vicuna-7B/13B Dai et al. (2023), Qwen-VL Bai et al. (2023), Qwen2-VL-Instruct-7B/72B Wang et al. (2024a), CogVLM-Chat Wang et al. (2024b), GPT-4o-mini and GPT-4o OpenAI (2024). These VLMs include both proprietary and open-source models, as well as various model sizes, which lead to variability in the generated target text. The details of prompts and hyperparameters used for each VLM are provided in Appendix C.2.
@@
Five expert annotators were provided with an explanation of the vision-language tasks and instructed to carefully review the definitions of the evaluation criteria. They then rated each target text for each criterion on a five-point scale. After completing the evaluations, they conducted a thorough review to ensure consistency. Each instance was scored independently by three annotators, resulting in a total of 18,000 human judgment scores.
@@
To collect overall judgments, we adopted a best-of-three approach, in which annotators are asked to choose the best text from three candidates. We did not use a five-point scale for the overall score because defining each rating level (from 1 to 5) for overall quality is difficult and could introduce bias Chiang et al. (2024).
Figure 4: Score distributions of the MMHE dataset.
@@ 3.3 Data Analysis
Figure 4 presents the score distributions for each criterion across the four tasks. The correctness and completeness criteria exhibit diverse score distributions for most tasks, suggesting that even the state-of-the-art VLMs face challenges in these aspects. In contrast, the fluency criterion shows a narrow score distribution across most tasks, with a dominant score of 5. This indicates that the VLMs generate fluent text even when visual understanding is inaccurate. Nonetheless, we consider fluency an essential criterion, as overlooking it could lead to high scores being assigned to correct but non-fluent text. A qualitative example illustrating this is provided in Section 4.3. The clarity and conciseness criteria have score distributions that are intermediate between the diverse and narrow distributions. While the score of 5 is more prevalent than in the diverse distributions, other scores still occur with noticeable frequency.
@@
HarmonicEval (ours) 66.6 76.4 73.4 77.0 73.4
Table 2: Overall accuracy (%) on MMHE. The best result for each task is marked in bold. Average (Avg.) indicates the average performance across the four tasks.
Metric REG VQA VDU IC
Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc
@@
Table 3: Criterion-wise correlations on MMHE. Scores for the most positively and negatively correlated criteria are marked with black and gray underlines, respectively. Crt: Correctness, Cmp: Completeness, Clr: Clarity, Flu: Fluency, Cnc: Conciseness.

### v2

# Multi-modal, Multi-task, Multi-criteria Automatic Evaluation with Vision Language Models

@@ Abstract
Vision-language models (VLMs) have shown impressive abilities across a range of multi-modal tasks. However, existing metrics for evaluating the quality of text generated by VLMs typically focus on an overall evaluation for a specific task, such as image captioning. While the overall evaluation is essential for any task, the criteria prioritized can differ depending on the task, making it challenging for current metrics to adapt to multi-task scenarios. To address this limitation, we propose HarmonicEval, a reference-free comprehensive evaluation metric that aggregates criterion-wise scores to produce the overall score in a bottom-up manner. Furthermore, we construct the Multi-task Multi-criteria Human Evaluation (MMHE) dataset, which comprises 18,000 expert human judgments across four multi-modal tasks. Our experiments demonstrate that HarmonicEval achieves higher correlations with human judgments than conventional metrics while providing numerical scores for each criterion [^1] Our code and data will be available publicly..
@@ 1 Introduction
Automatic evaluation of text generated by vision-language models (VLMs) is essential for improving their performance across various multi-modal tasks, such as image captioning and visual question answering Hessel et al. (2021); Lee et al. (2024b). As the range of tasks that VLMs can perform continues to expand, developing specialized evaluation metrics for each task becomes increasingly difficult. Hence, a comprehensive metric capable of evaluating text across multiple tasks is highly desirable. However, most existing metrics focus on measuring the overall quality of text within a specific task, limiting their applicability in multi-task settings (see Figure 1 (a)).
Figure 1: Multi-task and multi-criteria evaluation. (a) Conventional single-criterion approach focuses on a single task, such as image captioning. (b) HarmonicEval integrates multiple criteria to provide overall scores. MMHE dataset consists of 18,000 expert human judgments across four multi-modal tasks and five criteria.
Existing metrics that provide only overall scores Papineni et al. (2002); Zhang et al. (2020); Hessel et al. (2021) often prioritize specific evaluation criteria, as discussed in previous studies Kasai et al. (2022); Fabbri et al. (2021). For example, metrics for evaluating image captions typically prioritize correctness and completeness over conciseness and fluency. When these metrics are applied to other tasks, such as visual question answering, they tend to overvalue verbose or unnatural responses. To address this limitation, integrating multiple evaluation criteria to predict the overall score, a concept we refer to as comprehensive evaluation, holds significant potential for a more comprehensive assessment in multi-task scenarios. However, this approach remains underexplored due to the lack of meta-evaluation datasets that provide human judgment annotations across multiple tasks and criteria. This motivates us to introduce a novel evaluation metric with a new dataset in this paper.
@@
Figure 2: HarmonicEval framework consists of two steps. (a) Criterion-wise scoring is performed by prompting a VLM to evaluate the input text based on each criterion, followed by score smoothing to improve robustness based on the first-order statistics. (b) Score aggregation produces an overall score using harmonic weighting based on the second-order statistics, aiming to reduce statistical fluctuations.
First, we propose HarmonicEval, a reference-free harmonic evaluation metric for multiple multi-modal tasks. As shown in Figure 1 (b), HarmonicEval integrates multiple criteria to produce the overall score in a bottom-up manner. Specifically, the evaluation pipeline consists of two steps: 1) criterion-wise scoring, where a VLM is prompted to evaluate the input text based on each specific criterion, and 2) score aggregation, where the overall score is calculated from the criterion-wise scores. For score aggregation, we introduce a novel harmonic weighting scheme that automatically determines weight coefficients based on the second-order statistics of the output token probability distributions.
@@
Second, we introduce the multi-task multi-criteria human evaluation (MMHE) dataset, the first meta-evaluation dataset that provides human judgment annotations across multiple criteria and multiple multi-modal tasks. Specifically, MMHE consists of 18,000 expert human judgments on five criteria for four diverse tasks: referring expression generation (REG), visual question answering (VQA), visual document understanding (VDU), and image captioning (IC). This dataset serves as a benchmark for assessing the generalizability of automatic evaluation metrics.
Our experiments on MMHE show that HarmonicEval achieves higher correlations with human judgments than conventional metrics, while also providing criterion-specific scores that highlight areas for improvement. Furthermore, we demonstrate that HarmonicEval achieves state-of-the-art or comparable performance in conventional image caption evaluation scenarios across five widely used datasets that provide only overall judgments: Flickr8k-EX / CF Hodosh et al. (2013), Composite Aditya et al. (2018), PASCAL-50S Vedantam et al. (2015), and FOIL Shekhar et al. (2017). In summary, our key contributions are threefold:
  • We propose HarmonicEval, a novel reference-free metric for harmonic evaluation across multiple multi-modal tasks.
  • We introduce MMHE, the first multi-task multi-criteria human evaluation dataset, consisting of 18,000 expert judgments spanning four multi-modal tasks and five evaluation criteria.
  • We demonstrate the effectiveness of HarmonicEval on MMHE and five conventional image captioning datasets. In addition, we conduct the first in-depth analysis of how existing metrics implicitly prioritize different evaluation criteria.
2 HarmonicEval
As shown in Figure 2, the pipeline of HarmonicEval consists of two steps: criterion-wise scoring (§2.1) and score aggregation (§2.2).
2.1 Criterion-wise scoring
In this step, a VLM is employed as an evaluator and prompted to generate evaluation scores on each criterion independently. Let $\bm{t}$ be an input text to be evaluated, such as an image caption for the IC task. The evaluation process to obtain criterion-wise scores $s_{c}$ is formulated as $s_{c}=f([\bm{p}_{c},\bm{t}],\bm{x})$, where $c$ is a criterion, $\bm{x}$ is an input image, $f$ is a VLM, $\bm{p}_{c}$ is a prompt, and [,][,][ , ] denotes textual concatenation. To improve alignment with human judgements, score smoothing Liu et al. (2023b); Lee et al. (2024b) is applied as $\tilde{s}_{c}=\sum_{r\in R}rP(r\mid[\bm{p}_{c},\bm{t}],\bm{x})$ where $P(r\mid[\bm{p}_{c},\bm{t}],\bm{x})$ is the output token probability of for token $r$ assigned by the VLM, and $R$ is a set of ratings.
2.2 Score aggregation
To aggregate the criterion-wise scores $\tilde{s}_{c}$, we introduce harmonic weighting, a novel approach that leverages the second-order statistics of the output token probability distributions to adaptively determine the weight coefficients for aggregation. Compared to simple averaging, our aggregation approach aims to better align with human evaluation by dynamically emphasizing more reliable scores based on the input. Specifically, the overall score $S$ is computed as
$S=\sum_{c\in\mathcal{C}}w_{c}\tilde{s}_{c},\quad w_{c}=\frac{1}{H}\sigma_{c}^{-2(1-\gamma)/\gamma},$ (1)
where $w_{c}$ is a weight coefficient, $\mathcal{C}$ is a set of criteria, $\sigma_{c}$ is the standard deviation of the criterion-wise score given by
$\sigma_{c}=\sqrt{\sum_{r\in R}(r-\tilde{s}_{c})^{2}\ P(r\mid[\bm{p}_{c},\bm{t}],\bm{x})},$ (2)
and $H=\sum_{c\in\mathcal{C}}\sigma_{c}^{-2(1-\gamma)/\gamma}$ is the harmonic mean of the variances with a hyperparameter $\gamma$. Smaller values of $\sigma_{c}$ can be interpreted as indicating higher confidence in the evaluation of $c$. The role of hyperparameter $\gamma$ is to bridge three weighting strategies: uniform weighting, inverse variance weighting and selective weighting as detailed below.
Uniform weighting.
When $\gamma$$=$$1$, harmonic weighting reduces to uniform weighting $w_{c}$$=$$1/|\mathcal{C}|$. This is effective when all criterion-wise scores are equally reliable in determining the overall score. However, this does not provide the best estimator as observed variances are ignored in aggregation.
Inverse variance weighting.
When $\gamma$$=$$0.5$, harmonic weighting reduces to inverse variance weighting $w_{c}\propto\sigma_{c}^{-2}$. This provides the best linear unbiased estimator under the assumption that the observed variance is due to statistical fluctuations.
Selective weighting.
Discussion. When the evaluation criteria are carefully designed, the uniform weighting ($\gamma=1.0$) aligns closely with human expert judgment, and $1.00.5\leq\gamma<$ 1.0 further improves the alignment because it adaptively reflects the confidence of criterion-wise scores. Since the assumption underlying the inverse variance weighting ($\gamma=0.5$) is not reasonable when each criterion has its own variance, we hypothesize that a value between 0.5 and 1.0 is optimal and choose $\gamma=0.75$ as the default value. The detailed mathematical background is provided in Appendix A.
2.3 Implementation details
@@
Aggregating these five criteria contributes to the overall text quality evaluation across a wide range of tasks. Depending on the task and the style of the input text, some criteria may not be necessary. However, when such evaluations are conducted, the output becomes less confident, leading to higher variance $\sigma_{c}$ and lower weight coefficients $w_{c}$ as VLMs account for task and criterion features in addition to the input text. Thus, HarmonicEval can adaptively perform comprehensive evaluations without manual tuning of weight coefficients. More details on criteria selection are in Appendix B.
@@
Table 1: Five criteria for HarmonicEval and MMHE.
Prompts. The prompt $\bm{p}_{c}$ instructs the VLM to evaluate text with respect to the criterion $c$ on a five-point scale $R=\{1,2,3,4,5\}$. Below is the prompt of the correctness criterion for the IC task.
Your task is to rate the caption for the given image on a scale of 1 to 5 on the following criterion and rating scale.
@@
Here, the underlined portions indicate the task-dependent phrases. For example, “caption” is replaced with “answer” for the VQA task. The full prompts are provided in Appendix C.1.
3 MMHE Dataset
We present the MMHE dataset, the first meta-evaluation dataset that covers multiple evaluation criteria across multiple multi-modal tasks. We collected 18,000 expert human judgments spanning four tasks and five criteria. Example human judgment scores are shown in Figure 3.
3.1 Dataset design
Motivation.
Despite numerous human evaluation datasets for multi-modal tasks, most focus on a single task (e.g., image captioning) or offer only an overall quality label. The purpose of MMHE is twofold: 1) to assess how evaluation metrics perform in a multi-task setting, and 2) to verify the hypothesis that each task may prioritize different criteria (e.g., correctness, completeness). To achieve these goals, we design MMHE to cover multiple multi-modal tasks and to explicitly collect human judgments across five evaluation criteria.
Multi-modal tasks.
We select four diverse tasks to illustrate how different criteria might matter in different contexts: 1) REG aims to generate a textual expression that uniquely identifies a specific object in the image (marked by a bounding box). We expect that completeness and clarity are crucial for precisely distinguishing the target object. 2) VQA requires generating an answer to a question about the image content. Due to the nature of answering questions, we hypothesize that correctness and conciseness may be especially important here. 3) VDU focuses on interpreting and extracting information from visually presented documents. Similar to VQA, we suspect that correctness and conciseness play key roles. 4) IC involves producing a descriptive sentence for the entire image. We anticipate that correctness and completeness are especially relevant for capturing key elements.
By spanning these tasks within a single dataset, MMHE enables a fine-grained, criterion-wise analysis of how metrics handle varied requirements. It is worth noting that MMHE is the first dataset to unify multiple multi-modal tasks and multiple evaluation criteria in a single benchmark.
@@ 3.2 Dataset construction
The dataset construction process consists of three steps: 1) Source selection, which selects source text-image pairs for each task; 2) Target generation, which creates target texts to be evaluated using state-of-the-art VLMs; and 3) Human expert evaluation, which assesses the quality of the target texts.
@@
We selected the following four datasets: RefCOCO Kazemzadeh et al. (2014) for REG, OK-VQA Marino et al. (2019) for VQA, VisualMRC Tanaka et al. (2021) for VDU, and MSCOCO Lin et al. (2014) for IC. We randomly sampled 100 instances from the validation or test subset of each dataset.
@@
The target texts to be evaluated were generated using state-of-the-art VLMs. Specifically, we employed ten VLMs: LLaVA-1.5-7B/13B Liu et al. (2023a), InstructBLIP-Vicuna-7B/13B Dai et al. (2023), Qwen-VL Bai et al. (2023), Qwen2-VL-Instruct-7B/72B Wang et al. (2024a), CogVLM-Chat Wang et al. (2024b), GPT-4o-mini and GPT-4o OpenAI (2024). These VLMs include both proprietary and open-source models, as well as various model sizes, which lead to variability in the generated target texts. The details of prompts and hyperparameters used for each VLM are provided in Appendix C.2.
@@
Five expert annotators were provided with an explanation of the four multi-modal tasks and instructed to carefully review the definitions of the five evaluation criteria in Table 1. They then rated each target text for each criterion on a five-point scale. After completing the evaluations, they conducted a thorough review to ensure consistency. Each instance was scored independently by three annotators, resulting in a total of 18,000 human judgment scores.
@@
To collect overall judgments, we adopted a best-of-three approach, in which annotators are asked to choose the best text from three candidates. We did not use a five-point scale for the overall score because defining each rating level (from 1 to 5) for the overall quality is difficult and could introduce bias Chiang et al. (2024).
Figure 4: Human judgment score distributions for each task and criterion on the MMHE dataset.
@@ 3.3 Data analysis
Figure 4 presents the score distributions for each criterion across the four tasks. The correctness and completeness criteria exhibit diverse score distributions for most tasks, suggesting that even the state-of-the-art VLMs face challenges in these aspects. In contrast, the fluency criterion shows a narrow score distribution across most tasks, with a dominant score of five. This indicates that the VLMs generate fluent text even when visual understanding is inaccurate. Nonetheless, we consider fluency an essential criterion, as overlooking it could lead to high scores being assigned to correct but non-fluent text. The clarity and conciseness criteria have score distributions that are intermediate between the diverse and narrow distributions. While the score of five is more prevalent than in the diverse distributions, other scores still occur with noticeable frequency.
@@
GPT-FLEUR 60.1 75.0 76.5 76.0 71.9
HarmonicEval 66.6 76.4 73.4 77.0 73.4
Table 2: Accuracy (%) on MMHE. The best result for each task is marked in bold. Average (Avg.) indicates the average accuracy across the four tasks.
REG VQA VDU IC
Metric Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc Crt Cmp Clr Flu Cnc
@@
Table 3: Criterion-wise correlation analysis on MMHE. Scores for the most positively and negatively correlated criteria are marked with red and blue underlines, respectively. The highest correlations for each criterion are highlighted in bold. Crt: Correctness, Cmp: Completeness, Clr: Clarity, Flu: Fluency, Cnc: Conciseness.